OpenStudy (anonymous):

Use the Pythagorean Theorem to determine the distance between points A(5, 3) and B(-3, -3).

OpenStudy (anonymous):

First their distance away in the x axis is x2-x1 so 5-(-3) =8 Next in the y axis y2-y1 3-(-3)=6 (Important we used x2 and y2 from A because it's the more positive value, but if you had used from B it would come out as negative and when you square it it would still work, just be consistent, use x2 and y2 both from A or both from B) Next just use the pythagorean theorem and since you're looking for the hypotenuse, just calculate as: \[\sqrt{8^{2}+6^{2}}\]

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

Actually, I thought about it and you don't have to choose both from A or both from B. Just choose the bigger number of x as x2 and the smaller as x1 and same thing for the y. Bigger is y2 and smaller is y1.